# PsyDI: Towards a Personalized and Progressively In-depth Chatbot for Psychological Measurements

### v1

@@ 5.2 Score Model Evaluation
Table 1 shows the performance of PsyDI-en and PsyDI-ch on the benchmarks, compared with all the open-source and closed-course LLMs. As shown in the table, both PsyDI score models surpass all close-source models and open-source models on all the English benchmarks.
Specifically, on the Reddit and Reddit-mix datasets, the PsyDI models outperformed other models by at least 9.5%. On the Reddit-repeat dataset, many LLMs that had not undergone fine-tuning exhibited the issue of assuming that longer sentences contain more information, resulting in nearly half of the models achieving accuracies below 50%. In contrast, PsyDI was able to accurately identify completely redundant semantics 100% of the time and assign lower scores to repetitive long sentences. For the 16Personalities-EN dataset, due to its typical expressions, open-source models such as GPT-4, Qwen, and Baichuan achieved approximately 82% accuracy. However, PsyDI-en and PsyDI-zh achieved accuracies of 83.4% and 85.7%, respectively, slightly surpassing the best open-source models.
On the Chinese Benchmarks, Although PsyDI-en was not trained on Chinese datasets, its performance is still better than most models. After fine-tuning on Chinese datasets, PsyDI-zh has achieved results that surpass those of other LLM 7.5% on Diamonte. On the typical dataset 16Personalities-ZH, PsyDI-ZH is comparable with yi-medium. Moreover, it can be observed that GPT-4 performs worse than GPT-3.5 on some benchmarks. This is because GPT-4 chooses to say "I don’t know" for questions it is uncertain about, rather than providing an output.
Figure 4: Variations in scoring of score model with the inclusion of INFP expressions
To observe the scoring trends of the score model across four dimensions, we selected an original statement and gradually added expressions that are inclined towards the INFP four dimensions to the statement, observing the changes in the score model’s scoring. The results can be seen in Figure 4, where we present the score model’s scoring with different statements added, and highlight the target dimension. Note that the scores here have been multiplied by the label and normalized, so the closer the scores are to 1, the more likely they are to correspond to the relevant dimension. From the results, we can observe that by adding expressions of certain MBTI traits, the score model’s scoring does indeed gradually increase towards the corresponding dimension. At the same time, as more expressions of certain traits are added, the upward trend of the score model’s scoring in this dimension will also gradually slow down. This is because the score model is trained to evaluate the representativeness of sentences, so a sentence that is already very "INFP" will not have a significant change in representativeness even if more expressions of INFP are added. At the same time, we also observe that when adding expressions of certain traits, the non-target dimensions will also change. For example, after adding the statement that represents being alone, which is the I trait, "It’s in these solitary moments that I truly recharge and find my peace," the score for the N dimension also rises. This further illustrates that there are implicit associations between the four dimensions of MBTI.
@@
Figure 5: The score of sentences with extroverted semantics but introverted words
@@
Further, we wish to analyze whether the score model’s response to MBTI traits is at the word level or the semantic level. Therefore, we designed some sentences where the key words are of a certain MBTI trait, but the overall semantics are opposite to that trait. For example, in Figure 5, We started with an introverted sentence and replaced some of the words, such as verbs and predicates, with their antonyms. This resulted in a sentence where the key words remained introverted, but the overall semantics shifted to extroverted. We then employed our score model to evaluate these modified sentences along the I-dimension. The results showed that when the sentence’s meaning shifted, the score model accurately detected this change, assigning lower I-dimension scores to sentences with extroverted semantics. Detailed changes in scores across other dimensions are provided in Appendix E.1. This experiment underscores that the PsyDI score model is not merely responsive to individual keywords but is capable of comprehending and evaluating the full semantic context of sentences.
@@ 5.3 Ablation Study
In this section, we want to discuss all the effect of all the details in PsyDI. We discuss five five important design in PsyDI base on Reddit dataset:
@@
Table 2: Scoring models with varying classification heads and loss functions
@@
Prompt Design
Only post 69.2% 57.8% 24.8%
@@
Table 3: Scoring models with different training settings
Classification head. PsyDI apply a classification head number of four, corresponding to the four dimensions in MBTI, which are I/E, N/S, T/F, P/J. In principle, we hypothesize that different MBTI types require distinct scoring mechanisms, making a single head insufficient to capture the variations in scores across different MBTI types. Additionally, the 16 MBTI types are not entirely independent; their differences are proportional to the divergence in their underlying dimensions. For instance, the difference between INFP and ENFP is relatively smaller than the difference between INFP and ESTJ. To address this, we designed the output layer of the scoring model in PsyDI with 4 heads.
In this study, we investigate the impact of varying the number of heads on the performance of our model using the Reddit dataset in Table 5.3. In the one-head setup, a single head directly predicts the MBTI value for the given input. The performance across three datasets was significantly low, with an accuracy of only 5% on the mixed dataset. This poor performance can be attributed to the ambiguity introduced by having a single head predict all MBTI types, making it challenging for the model to determine the representativeness of each MBTI type accurately.
In the 16-heads setup, each MBTI type has a separate output head. While evaluation uses the corresponding head, performance remains low across all datasets. This is due to data being split into 16 parts, leading to insufficient training for each head and the inability to capture interdependencies among MBTI dimensions.
Loss function. The objective of the score model is to determine the ranking of pair-wise data. There are several loss functions that can achieve this objective. In Table 5.3, we compare the performance of different loss functions. The pair-wise loss is defined as follows:
$L(p_{i},p_{j},mbti)=-\log(\sigma(mbti^{\top}\cdot(\mathcal{F}(p_{i})-\mathcal{F}(p_{j}))))$ (7)
The multi-margin loss is defined as follows:
$L(p_{i},p_{j},mbti)=\max(0,margin-mbti^{\top}\cdot(\mathcal{F}(p_{j})-\mathcal{F}(p_{i})))$ (8)
Both of these loss functions performed significantly worse than the existing score model across three datasets.
Prompt. The prompt for PsyDI’s score model is designed as:
Please predict to what extent the posts can be assessed as <MBTI> for this user.
To validate the necessity of this prompt design, we conducted a comparison of scoring with and without the prompt on the Reddit dataset. The result is presented in Table 5.3As shown, the accuracy on the mixed and repeated datasets is lower compared to the performance under the complete prompt. This discrepancy arises from the model’s difficulty in distinguishing which classification task to perform when not provided with the specific MBTI prompt. For instance, in the mixed dataset, $p_{i}$ might be an INFP statement, while $p_{j}$ could consist of 30% INFP and 70% ESTJ statements. Without the explicit prompt, the model may incorrectly interpret the task as identifying the representativeness of ESTJ, thus assigning a higher score to the latter. This highlights the importance of the complete prompt in guiding the model to the correct classification task.
Data Augmentation To avoid the influence of natural biases in Reddit data on the score model, such as longer sentences naturally containing more MBTI-related information than shorter ones, we performed data augmentation on the Reddit dataset by constructing mixed and meaningless datasets. Here, we discuss the need for these two data enhancements. We compared the differences between not using any data enhancement, using only mixed data, using only repeat data, and using the full score model. As shown in Table 5.3, under the joint training of the mixed dataset and the repeated dataset, not only do the performance metrics improve on the respective mixed and repeated test sets, but there is also a notable enhancement in performance on the Reddit dataset itself.
@@ 5.4 Pipeline Evaluation
Figure 6: scores of 16-type MBTI bot across all MBTI types during the PsyDI testing Process
To verify the effectiveness of the PsyDI pipeline, we utilized the AI Character’s MBTI bot  [^3] https://character.ai/ to interact with PsyDI and observe its performance. Specifically, the whole test process of PsyDI involves 4-5 rounds of multi-turn QA, during which the MBTI profile is updated after each round. We monitored the changes in the MBTI profile following each QA session to assess the trend and accuracy of the MBTI profile updates. The result is shown in Figure 6, where each row represents for an MBTI bot, while each column represents for the MBTI score in MBTI profile.
As shown in the result, in the initial stage, when only the user’s statements are available, there is no significant correlation between PsyDI’s predicted MBTI and the actual MBTI. At this point, most of the MBTI scores are relatively low. By the third round of QA (stage 3), a clear correlation begins to emerge between PsyDI’s predicted MBTI and the bot’s actual MBTI. At this stage, the scores for the true MBTI and several closely related MBTIs are significantly higher than those for other MBTIs. By the final stage (stage 5), after 4-5 rounds of multi-turn dialogue, the scores for the true MBTI and similar MBTIs are much higher than for other MBTIs, indicating that PsyDI can effectively differentiate between different MBTI types. Additionally, since the MBTI bot’s expressions do not perfectly align with its actual MBTI, PsyDI also assigns relatively high scores to some similar MBTIs. From these scores, we can also observe which MBTI types are more alike; for instance, the scores for XNFX types tend to be close to each other.
5.5 Online Evaluation
Additionally, since PsyDI went live  [^4] The website address will be published after the review, it has been tested by over 1,000 users. Based on this substantial dataset, we conducted a thorough analysis of the collected data.
First, we analyzed the most commonly used words in the responses of users from each MBTI type, including their free-text answers, to observe the word frequency patterns for different MBTI types. We created word clouds for each dimension, such as for the I dimension, where we collected responses from all IXXX users and excluded words that frequently appeared in EXXX responses. This approach helped us isolate words specifically associated with the I/E dimension. The result for the I/E and F/T dimensions is shown in Figure 7 (a), while the remaining dimensions are presented in Appendix E.2.
Our findings indicate a strong correlation between the MBTI detected by PsyDI and the common word frequencies in users’ self-reports. For instance, users with the I dimension often used introspective words like "inner," while those with the E dimension frequently used socially interactive words like "help" and "share." Similarly, users with the T dimension tended to focus on problem-solving, whereas those with the F dimension were more concerned with helping and supporting others. These patterns align well with the expected behaviors of these MBTI dimensions, suggesting that PsyDI’s MBTI assessments are closely related to the word frequencies in users’ responses.
Additionally, we aimed to compare the patterns displayed by MBTI users in multimodal tasks. We selected a blob tree task from PsyDI, where users chose the blob they felt most resembled themselves. We calculated the proportion of users from each MBTI type who selected each blob and highlighted the top four significant positions in Figure 7 (b). For example, position 20, described as "ambitious, confident, and full of life," was predominantly chosen by EXTX users, reflecting their extroverted and logical characteristics. Position 15, chosen mainly by XNFX users, corresponded to openness to new experiences and friendliness, matching N and F tendencies. Position 10, associated with "hardworking and determined," was primarily chosen by XXTJ users, aligning with their leadership traits.
Thus, PsyDI’s MBTI assessments correspond well with the blob tree results, a type of projective test, further validating PsyDI’s effectiveness.
@@ 6 Related Works
6.1 Psychological Assessment
As one of the primary measuring methods in psychology, assessment has been developed a long history and has been established a comprehension measurement system covering different sub-domains, such as clinical diagnosis, military mental state assessment, educational improvement, business talent management and so on (Cohen et al., 1996). Psychological assessment has many categories, including testing, interview, behavior observation experiment, etc. Testing has been the most preferable tools in many scenery due to its preciseness based on well-designed criterion, consistent measurement despite scenery and convenience in re-usability. However, it needs sufficient expert opinions to design as well as lacks personalized testing to deepen the assessment (Merry et al., 2012). Moreover, expert interview make up these short comes, yet suffers from plenty of resources to collect sparse, non-mass-production-able instances (Crisp & Lincoln, 2014).
The Myers-briggs type indicator (MBTI) is a self-report questionnaire on personality types in personnel management (Myers et al., 1962). This test and current concepts are originated from Carl Jung’s personality types, a typology theory building and investigating an individual’s mental cognitive model (Jung, 1923) and has been extended into 16 personalities by John Beebe (Jung & Beebe, 2016). Although originated from but not belongs to psychological research field, the assessment method of MBTI test has been discussed as having reference value and also gain a wide accept between many individuals (Boyle, 1995; Lamond, 2001).
@@ 6.2 Large Language Models towards Psychological and MBTI
In the advantage of capability emergence firstly evolved out by OpenAI’s GPT models, large language models could spontaneously capture the deep understanding in contextual relevant contents, achieving human-like level performance (Achiam et al., 2023b). Various works have discussed the pros and cons large language models has reached in psychological research field, demonstrate the potential capabilities LLMs mastered in psychological chatting, analysing, guiding and reasoning (Ke et al., 2024; Demszky et al., 2023; Binz & Schulz, 2023). Current cross-over applications includes multi-turn dialogue framework for psychological counseling Zhang et al. (2024b), conversational diagnose (Tu et al., 2024b), gamified assessments (Yang et al., 2024), introducing more noval pipelines under this scope.
Also been a growing force in psycho-cognitive research, LLMs have shown a capable solution to recognize and simulate human cognitive process (Bubeck et al., 2023). This helps MBTI-related AI application exploration whose essence is human cognitive patterns. A series of works have investigated how LLMs captures cognitive personaliies via role-playing different types (Chen et al., 2024; Cui et al., 2023; Tseng et al., 2024; La Cava et al., 2024; Pan & Zeng, 2023). LLM-driven assessment abilities are studied to confirm the MBTI awareness also contributes personality discrimination (Rao et al., 2023; Song et al., 2024). A gamified MBTI test has been designed and evaluated on PsychoGAT (Yang et al., 2024).
@@ 7 Conclusions and Limitations
In this paper, we delve into the comparative analysis of psychological assessment AI with traditional methodologies like scales, questionnaires, and expert interviews. We conceptualize the multi-turn dialogues between the LLM agent and the test taker as a MDP and Goal-Conditioned RL problem. However, It is crucial to recognize the challenges associated with directly amassing extensive data of multi-turn dialogue states and rewards. Instead of pursuing a end-to-end RL agent, we draw some insights from psychological scales and propose a unified framework named PsyDI. The framework mainly consists of a pipeline and a score model. The former utilizes the ability of LLMs to adaptively generate engaging multi-turn question-answers and designs a MBTI profile to monitor and quantify the MBTI type of the taker at different conversation stages. The latter incorporates several data-generation methods like masked rank pair, and network optimization techniques such as ranking loss and 4-heads prediction. Our quantization experiment results and visualization analysis about the public verification version of PsyDI demonstrate its effectiveness and potential for interactive psychological assessments. Also, it is noteworthy that current PsyDI is a multiple components prototype. Utilizing the sequential data and user feedback collected through PsyDI to train an end-to-end network model with RLHF algorithms presents a promising avenue for future exploration. Besides, it is also interesting to transform PsyDI from a turn-based chatbot into a streaming AI assistant or therapist, akin to face-to-face human interaction. This paradigm shift would incorporate a richer array of information, such as facial expressions and intonations, thereby enhancing the comprehensiveness of psychological evaluations.
Appendix A Dataset Details

### v3

@@ 5.2 Score Model Evaluation
Table 1 presents the comparative accuracy of Score Model in PsyDI against both open-source and closed-source models. Notably, PsyDI models consistently outperform all other models on both English benchmarks and Chinese benchmarks. Specifically, on English benchmarks, both PsyDI-EN and PsyDI-ZH lead by at least 9.5% on the Reddit and Reddit-mix datasets. On the Reddit-repeat dataset, where many un-finetuned LLMs erroneously prioritize longer sentences, PsyDI achieves near perfection in detectingsemantic redundancy, assigning lower scores to repetitive long sentences. On the 16Personalities-EN dataset, known for its distinctive expressions, PsyDI-EN and PsyDI-ZH achieve accuracies of 83.4% and 85.7%, surpassing open-source models which average around 82%. On Chinese benchmarks, PsyDI-EN, despite not being trained on Chinese data, still outperforms most models. Statement-fine-tuning on Chinese datasets, PsyDI-ZH surpasses other LLMs by 7.5% on Diamante.
To analyze the scoring changes of the model across the four MBTI dimensions, we conduct an experiment where a statement was augmented with phrases progressively aligned with the INFP traits. The evolution of the model’s scores is tracked as depicted in Figure 4. The scores are normalized to facilitate comparison, with values closer to 1 indicating a stronger alignment with INFP. The findings indicate that the model’s scores do incrementally favor the corresponding dimension as trait-specific expressions are appended. The rate of score increase slows down as the sentence becomes overly saturated with INFP traits, which is a result of the model’s training to assess the indicativeness of sentences. Concurrently, the addition of certain trait expressions induces changes in non-target dimensions, exemplified by the rise in the N dimension score when the introverted (I) trait expression "It’s in these solitary moments that I truly recharge and find my peace" is included. This further illustrates that there are implicit associations between the four dimensions of MBTI.
@@
Figure 5: The score of sentences with extroverted (E) semantics but introverted (I) words
@@
Also, we conducted an experiment to determine whether the PsyDI score model assesses MBTI at the word level or the semantic level. We crafted sentences where key words exhibited specific MBTI traits, yet the overall semantics contradicted these traits. In Figure 5, we initiated with an introverted sentence and substituted certain words, including verbs and predicates, with their antonyms. Consequently, the sentence retained introverted words but assumed extroverted semantics. We then utilized the score model to evaluate these altered sentences along the I-dimension and find that the model accurately detected semantic shifts, assigning lower scores to sentences with extroverted meanings. More score variations are detailed in Appendix E.1. This experiment confirms that the score model is adept at interpreting and evaluating the complete semantic context of sentences, rather than merely reacting to individual keywords.
@@ 5.3 Ablation Study
Table 2: Ablations about score model variants with different classification heads and loss functions.
@@
MBTI-goal Prompt
w/o MBTI-goal 69.2% 57.8% 24.8%
@@
Table 3: Ablations about score model variants with different prompt and data settings.
In this section, we aim to elucidate the impacts of the important designs of PsyDI on Reddit dataset, specifically the classification head, loss function, mbti-goal prompt, and data augmentation. The results depicted in Table 5.3 and  5.3 demonstrate that the substitution of these elements leads to a noticeable decline in model performance, underscoring their necessity. For detailed explanations, please refer to Appendix E.3.
@@ 5.4 Pipeline Evaluation
Figure 6: INFP scores with random initializations of MBTI profile.
Figure 7: MBTI rank of 16 MBTI bots under three
distinct MBTI Testing.
@@
Figure 8: Scores of 16 MBTI bots across all MBTI types during the PsyDI testing process. Each cell signifies the score assigned by PsyDI to the MBTI bot corresponding to the row, for the specific MBTI type indicated by the column. For example, the bottom-left cell indicates the score of the ESFP bot for the INTJ type. Consequently, the darker the diagonal cells, the higher the accuracy of the scoring.
In this section, we assess the effectiveness of the PsyDI pipeline from three perspectives. First, given the variability and potential inconsistencies in these self-reports, it is imperative for PsyDI to iteratively refine the user’s MBTI profile through continuous updates, irrespective of the initial statements’ reliability. Thus, in Figure 7, we initiate the experiment with several randomized MBTI profiles and utilize them for further assessment with an INFP bot provided from Character AI (CharacterAI, 2024). Initially yielding INFP prediction scores between 35 and 50, the scores consistently rise and eventually stabilize above 80. This indicates that PsyDI effectively refines the MBTI profile through extensive dialogue in the pipeline, even with an unfavorable initial state.
Following this, we perform a detailed comparison between PsyDI and several prominent MBTI tests regarding their rank precision in identifying MBTI. 16Personalities and APESK (APESK, 2024) are two extensively acknowledged MBTI tests. We utilize the 16 most commonly used MBTI bots from Character AI and have each bot respond to the questions from these two tests and PsyDI. We document the rank of the bot’s actual MBTI type among all other MBTI types under a specific MBTI test. The comparative results are depicted in Figure 7, where the horizontal axis represents different MBTI bots, with each point indicating the rank of the bot’s actual MBTI. The results reveal that PsyDI’s average ranking accuracy is comparable or superior to that of the two established tests. Moreover, for the majority of bots, their actual MBTI type is among the top three predictions made by PsyDI, which demonstrates greater stability compared to others.
Thirdly, we also monitor the test interaction procedure of the Character AI bot with PsyDI. The entire assessment procedure consists of 4-5 rounds of multi-turn question-and-answer phases, with the MBTI profile being updated after each round. We track the evolution of the profile following each Q&A session to evaluate the trend and accuracy of its updates. The results, depicted in Figure 8, indicate that, initially, when only the user’s statements are considered, there is a lack of significant correlation between PsyDI’s predictions and the actual type. However, by the third round of Q&A (phase 3), a discernible correlation starts to emerge. By the concluding stage (phase 4), the scores for the true MBTI and the prediction are substantially higher than those for other types, demonstrating PsyDI’s capability to accurately distinguish between various types. Furthermore, due to the bot’s expressions not perfectly reflecting its actual MBTI, PsyDI also assigns higher scores to some similar MBTIs, thereby revealing which types are more closely related; for example, scores for XNFX types, represented by the 4x4 matrix in the top-left corner, tend to cluster together.
In Appendix E, we introduce supplementary experiments, encompassing a semantic scoring experiment aimed at discerning the scoring tendencies of the score model in Appendix E.1, along with an exhaustive analysis of the data collected from PsyDI Online in Appendix E.2. Moreover, we have extended the PsyDI pipeline to encompass additional psychological topics to demonstrate its scalability, as shown in Appendix E.4.
@@ 6 Related Works
6.1 Psychological measurement
As one of the primary measuring methods in psychology, measurement has been developed a long history and has been established a comprehension measurement system covering different sub-domains, such as clinical diagnosis, military mental state measurement and so on (Cohen et al., 1996). Psychological measurement has many categories, including testing, interview, behavior observation experiment, etc. Testing has been the most preferable tools in many scenery due to its preciseness based on well-designed criterion, consistent measurement despite scenery and convenience in re-usability. However, it needs sufficient expert opinions to design as well as lacks personalized testing to deepen the measurement (Merry et al., 2012). Moreover, expert interview make up these short comes, yet suffers from plenty of resources to collect sparse, non-mass-production-able instances (Crisp & Lincoln, 2014). Although originated from but not belongs to psychological research field, the measurement method of MBTI test has been discussed as having reference value and also gain a wide accept between many individuals (Boyle, 1995; Lamond, 2001).
@@ 6.2 Large Language Models towards Psychological and MBTI
In the advantage of capability emergence firstly evolved out by OpenAI’s GPT models, LLMs could spontaneously capture the deep understanding in contextual relevant contents, achieving human-like level performance (Achiam et al., 2023). Various works have discussed the pros and cons large language models has reached in psychological research field, demonstrate the potential capabilities LLMs mastered in psychological chatting, analysing, guiding and reasoning (Ke et al., 2024; Demszky et al., 2023; Binz & Schulz, 2023). Current cross-over applications includes multi-turn dialogue framework for psychological counseling Zhang et al. (2024), conversational diagnose (Tu et al., 2024), gamified assessments (Yang et al., 2024), introducing more noval pipelines under this scope. Also been a growing force in psycho-cognitive research, LLMs have shown a capable solution to recognize and simulate human cognitive process (Bubeck et al., 2023). This helps MBTI-related AI application exploration whose essence is human cognitive patterns. A series of works have investigated how LLMs captures cognitive personaliies via role-playing different types (Chen et al., 2024; Cui et al., 2023; Tseng et al., 2024; La Cava et al., 2024). LLM-driven measurement abilities are studied to confirm the MBTI awareness also contributes personality discrimination (Rao et al., 2023; Song et al., 2024). A gamified MBTI test has been designed and evaluated on PsychoGAT (Yang et al., 2024).
@@ 7 Conclusions and Limitations
In this paper, we delve into the comparison of psychological measurement AI with traditional methods like scales and expert interviews. We conceptualize the multi-turn dialogues between the LLM agent and the test-taker as a MDP problem. However, It is crucial to recognize the challenges associated with amassing extensive data of multi-turn dialogues. Instead of pursuing an end-to-end RL agent, we draw some insights from psychological scales and propose a unified framework named PsyDI, which consists of a pipeline and a score model. The former utilizes the ability of LLMs to adaptively generate engaging multi-turn Q&A and designs an MBTI profile to quantify user’s the MBTI. The latter incorporates several data-generation methods like masked rank pairs, and network optimization techniques such as ranking loss and 4-heads prediction. Our experiment results and visualization analysis about the public verification version of PsyDI demonstrate its effectiveness and potential for psychological interactions. Also, it is noteworthy that current PsyDI is a multiple components prototype. Utilizing the sequential data and user feedback collected through PsyDI to train an end-to-end network model with RLHF algorithms presents a promising avenue for future exploration. Besides, it is also interesting to transform PsyDI from a turn-based chatbot into a streaming AI therapist, akin to face-to-face human interaction. This paradigm shift would incorporate a richer array of information, such as facial expressions, thereby enhancing the comprehensiveness of psychological evaluations.
Acknowledgments
We thank several members in Shanghai AI Laboratory for their invaluable help and support, which have significantly enhanced the quality of this work. We are thankful to Qingzi Zhu, whose exceptional skills in graphic design were pivotal in creating the visual materials for both the online presentation and this manuscript. Additionally, we acknowledge Mengjie Jiang for her contributions in not only statistically analyzing the experimental data and crafting the key overview figure but also generalizing our framework into another psychology field. Their collective contributions have been indispensable, and we are deeply appreciative of their support.
Appendix A The Background of MBTI
The unique combination of these four dimensions results in sixteen distinct personality types, each with its own cognitive patterns and behavioral tendencies. This heterogeneity has contributed to the growing popularity of the MBTI in recent years, with over 1.207 billion test administrations recorded by 16Personalities. However, it is imperative to acknowledge that the MBTI is based on Carl Jung’s theory of psychological types (Jung, 1923). Jung’s theory postulated a quaternary of cognitive functions—thinking, feeling, sensation, and intuition—each with an orientation of either extroversion or introversion, resulting in eight dominant functions, Te (Extroverted Thinking), Ti (Introverted Thinking), Fe (Extroverted Feeling), Fi (Introverted Feeling), Se (Extroverted Sensing), Si (Introverted Sensing), Ne (Extroverted Intuition), and Ni (Introverted Intuition). The constitution of an individual’s personality is delineated by the hierarchical preference for these dominant functions. According to the theory of Jung, it is improbable for a person to exhibit high predilections for both extraversion and introversion within the same cognitive function, as human cognitive resources are finite. This principle leads to the emergence of sixteen common cognitive function sequences, aligning congruently with the sixteen MBTI types. For example, the cognitive function sequence of INFP is: Fi, Ne, Si, Te, Fe, Ni, Se, Ti. Consequently, the adjudication of MBTI typologies inherently entails the scrutiny of one’s cognitive processes. The cognitive preferences, on the other hand, refer to the individual’s preferred or dominant use of these functions.
Appendix B Experimental Settings
B.1 Datasets
We constructed the following MBTI statement datasets with labels in both English and Chinese:
  • Reddit. The Myers Briggs Personality Tags on Reddit Data (Storey, 2018) is a public dataset consisting of user statements and their self-reported MBTI types. To balance the distribution of MBTI types, we extracted 400 statements for each MBTI type, ensuring that each statement was longer than 100 words. The Reddit dataset is augmented with the data augmentation methods introduced in section 4.3, forming Reddit-mix and Reddit-repeat.
  • 16Personalities. We obtained all the questions from 16 Personalities  [^1] https://www.16personalities.com/, an authoritative MBTI testing site. Subsequently, we used one LLM to simulate specific MBTI types and engaged another LLM in multi-turn Q&A sessions with the former one. If the predicting LLM correctly predicted the MBTI type of simulating LLM, the Q&A record was collected along with the MBTI labels. We collect the dataset in both English and Chinese to form 16Personalities-en and 16Personalities-zh.
  • Diamente The Diamente Chinese chit-chat dataset comprises high-quality chit-chat conversations collected with the assistance of a pretrained dialogue model that aligns well with human values. Using these chit-chats as a starting point, we asked LLM to generate multi-turn Q&A sessions. We randomly selected answers and recorded the entire process, labeling the statements based on LLM’s predictions to form dataset Diamente.